# Novel SiC UV Instrumentation Development with Potential Applications for the Habitable Worlds Observatory

### v1

@@ 3 Overview of development activities
SiC instrument development by our team under the current programs have leveraged the expertise of scientists, engineers and commercial companies with fabrication, integration and implementation expertise. Key to the fabrication tasks across all the project is the work that CoolCAD Electronics, LLC, has been carrying using unique expertise in the design and fabrication of devices and integrated circuits with SiC. CoolCAD has previously worked on SiC low-power CMOS, deep-UV sensors, and building blocks for deep-UV-sensitive focal plane arrays as well as high-power, high-voltage and high-temperature SiC devices. CoolCAD’s current work to obtain UV sensors and readout circuitry on the same substrate leverages background work supported by NASA SBIR programs (NNX17CG32P, “Silicon Carbide 10µm-Pitch UV Imaging Array and APD with Active Pixel Readout,” Phase 1 SBIR; NNX16CG51P, “Single Chip EUV, VUV, and Deep UV Photodetector System with Integrated Amplifier,” Phase 1 SBIR; NNX17CG15C, “A Silicon Carbide Foundry for NASA’s UV and High Temperature CMOS Electronics Needs,” Phase 2/2E SBIR). CoolCAD has fabricated/demonstrated two classes of UV sensor devices and structures, both of which are top-lit: PN-junction and Schottky diodes and SiC UV Focal Plane Arrays, which are the focus of the current projects. The spectral responsivity is partially dependent on depth of the photocurrent-collecting junction electric field. Schottky diodes, with the built-in field closest to the surface, are expected to show a responsivity peak with a shorter wavelength, which has been observed in the devices. CoolCAD employs implant profile engineering, to set the implant energies and doses necessary to achieve a desired profile, which has also been incorporated in current work. Finally, both past GOES-R funding and a current IRAD has supported demonstration of the radiation hardness and visible blind nature of the SiC material that is relevant to all of the projects under development.
@@
The PReSSiC project aims to deliver a lab-tested 1×128 pixel 4H-SiC detector suitable for spectrometers in miniaturized platforms that can obtain radiance observations with a higher detectivity than previous generations of NUV spectrometers from 200-340nm at a Rsimilar-$\sim$200. Key development for this project has been testing a number of design trades and optimizing design and fabrication for this goal. The project has begun with a focus on design and layout of individual sensors and small arrays. Potential single-pixel geometries and fabrication parameters are being evaluated to optimize device characteristics such as responsivity and response speed, which will be examined in the context of the science requirements. A key area of exploration is a test of implant depth, as it is valuable to create an implant profile which supports enhanced sensitivity in the spectral region of interest. There are limits on this process, such as difficulty of getting implants past a certain depth into the tough SiC crystal material. An additional consideration is that in a focal plane array, the pixels must be electrically and optically isolated for clear image formation - deep isolation trenches. CoolCAD Electronics has previously demonstrated small-pixel-count 2-D arrays with 10 µm pitch, that had not featured deep trench isolation, but is examining deep isolation trenches in the new arrays.
@@
The project is laying out and fabricating a number of different test structures and small-pixel-count arrays to identify potential fabrication difficulties and limits to our fabrication techniques. Diode pitch will be varied between 6-40 µm and the project is looking at a number of factors including: contact resistance, minimum trench width for effective fill and planarization, patterning long metal Lines, surface leakage reduction, and effects on responsivity and dark current. The team has also started examining packaging and I/O structure design for both active and passive pixels. The team is designing and building a set of readout electronics, particularly for planned scaling up of the pixel count for the full array. Radiometric testing is being conducted and will be conducted at numerous steps along the way and a testbed is currently being assembled so that the arrays can eventually be tested under both ambient and vacuum conditions. Key performance testing related to responsivity, noise performance and dispersion will be compared to requirements based upon simulations.
@@
Another key ongoing development effort is the NASA Early Career Initiative-supported “Design of a Broad Utility, Visible-Blind, SiC, UV Spectrometer for Astrophysics, Planetary and Earth Science Applications, Validated Through Representative Measurement of Climate Forcing Gases” project, hereafter referred to as the Pandora-SiC effort, which aims to achieve clean differential optical absorption spectroscopy measurements of columnar trace gases in the earth atmosphere in the 280-450 nm regime. This is accomplished through integration of visible blind UV detectors, produced leveraging recent advances in SiC fabrication, into an existing, ground-based, atmospheric spectroscopy platform, Pandora [11], which has been developed by a commercial partner, SciGlob, LLC. The ECI Team is hoping to improve the reliability and temporal resolution of total column formaldehyde measurements conducted by Pandora either for direct data collection or validation for data collected from satellite platforms like the ozone monitoring instrument (OMI) [24]. A simultaneous goal of this project is also to simplify instrument setup through elimination of filter components, enabled by the visible blindness of SiC, and reducing or eliminating cooling requirements, enabled by the large bandgap and therefore reduced dark current of SiC. The ECI effort is at the point of integrating fabricated SiC detectors, $1\times 128$ and $1\times 1024$ arrays of high aspect ratio ($25~{}\mu\text{m}\times 200~{}\mu\text{m}$, $25~{}\mu\text{m}\times 500~{}\mu\text{m}$, and $50~{}\mu\text{m}\times 500~{}\mu\text{m}$), passive, P-on-N photodiodes with external capacitive transimpedance amplifier (CTIA) read-out integrated circuits (ROIC) configured to be controlled by an external imaging array controller. Following integration, detector units will be tested to confirm quantum efficiency (QE) greater than 0.1 at 345 nm and dark currents $<20~{}\text{pA}$ (individual diodes probed so far show a dark current upper limit of $<6~{}\text{fA}$, discussed in section 5), and integrated into the Pandora spectrometer.
@@
The current detector and electronics configuration for both the PReSSiC detectors and the Pandora-SiC detectors has been tailored for easy integration into planetary science instruments and Pandora instrument systems, respectively. While not directly applicable to HWO, testing of pixel design, properties and layout in the PReSSiC project and maturation of the integration scheme developed through the ECI effort both provide key techology and engineering advancement for potential HWO detectors. Achieving desired noise characteristics, responsivity, and validating operation of SiC detectors using standard read out techniques offers a compelling platform for operation of denser and higher dimensionality SiC photodiode arrays of interest for use in potential Coronagraph, Spectrograph, and High Resolution Imaging Instruments.
@@ 4 Fabrication and Integration Progress of Photodetectors
For the PReSSiC and ECI projects, many of the initial tasks are focused on design and fabrication, and CoolCAD is working on a variety of linear photodiode structures in order to map the design space, identify design best practices, and optimize fabrication techniques.
Figure 1: Micrographs from the process of different diode regions being defined by ion implantation for the tall, narrow aspect ratio diodes being developed for spectroscopy applications for the ECI program.
Diode geometry is the first aspect being explored. For the particular spectroscopy application targeted by the ECI program, the pixel pitch required for the target spectral resolution is 25 $\mu$m. The optical design of the instrument can take advantage of pixels with high aspect ratio. Therefore the initial fabrication run of the program includes pixels with 25 $\mu$m x 500 $\mu$m and 25 $\mu$m x 200 $\mu$m drawn areas, along with test structures with a different aspect ratio. Figure 1 shows micrographs from the fabrication process of these designs, defining different diode regions. In the context of the PReSSiC program, a major program target is miniaturization to a very small pixel pitch and the devices being fabricated for this program include diodes with pitches as small as 10 $\mu$m and 6 µm. Figure 2 shows fabrication micrographs of some of these structures. Other geometric features being explored are the placement and sizes of contacts and interconnects, to evaluate certain trade-off aspects such as fill factor vs. diode series resistance.
@@
Another aspect is different junction structures. We are fabricating pn-junction diodes with different doping profiles for these two programs, which will yield different geometries and depths for the photocurrent collection regions. For a separate NASA SBIR program, we are also fabricating Schottky diodes, which have the photocurrent collection region starting at the illuminated surface. Since shorter wavelengths have larger absorption coefficients and smaller absorption depths, the different doping profiles are expected to have somewhat different spectral responsivity. Planned further developments include the use of different epitaxial structures for diode fabrication as well.
Figure 3: The wafer-scale layouts for the PReSSiC project fabrication run (left) and the ECI program fabrication run (middle, right).
For applications which use diode arrays, such as spectroscopy or imaging, pixel-to-pixel isolation can be required. For the ECI program, deep trenches between the photoactive regions of each diode are being implemented for this purpose. For the PReSSiC program, in addition to trenches, we are also implementing array variations using a specific diode geometry which was first tested for this purpose in another development program. This alternative approach allows us to increase the fill factor, especially of the smallest diodes. The development and evaluation of these isolation structures are being pursued in the SBIR program as well.
Figure 3 shows the wafer-scale designs for the initial runs in the PReSSiC project and the ECI program. The PReSSiC program design includes small-pixel-count arrays (15x1 and 8x1) and single test diodes. The ECI program design includes 128x1 and 1024x1 arrays and single test diodes. Both designs are currently in fabrication and have been through their implant schedule and isolation trench formation. We are working on the back-end processing, which comprises the metallization steps for contact metal and interconnect formation.
@@
For both projects, the next stage of development will include active diode arrays, for which the first stage of a readout circuit converting the photocurrent to a voltage level over an integration period is fabricated on the same die with the photodiode arrays. CoolCAD Electronics has active development programs on SiC MOSFET structures and circuits and has previously demonstrated single diodes integrated with this readout circuit. Figure 4 shows previously-fabricated examples of this structure. These development programs all start from the level of physics-based semiconductor device modeling and simulations, and include process simulation and development efforts as well as device and circuit design, layout, and fabrication. For the PReSSiC project, in addition to one-dimensional arrays for spectroscopy, two-dimensional arrays will also be implemented for push-broom spectroscopy and imaging applications. The pixel count will be expanded as well. In the SBIR program, CoolCAD is further developing larger scale two-dimensional arrays, both of passive and active pixels. For the two-dimensional active array development, this effort includes work to miniaturize the transistors as well to be able to integrate the circuit for each pixel within the pixel, since scaling a 2-D array to significant pixel counts can be impractical otherwise.
@@
Figure 5: The structure of SiC photodiode-based focal plane arrays for the ECI effort consists of: (a) A core 1×1024 array of SiC photodiodes, which fans out to wirebond pads matched to the inputs of wire bondable ROICs. In this case, the array is interfaced with 4 CTIA ROICs, each with 256 inputs. (b) The ECI effort is currently in the process of characterizing multiple diode variants, including: (i) arrays with a 25 µm pitch, and (ii) arrays with a 50 µm pitch.
The ECI effort is focused on a short-term field demonstration of SiC focal plane arrays in an existing atmospheric spectroscopy system, Pandora, for formaldehyde detection. As a result, the selection of diode array formats was informed by the need for a mature and highly reliable diode design, the ability to interface with external readout circuitry without complex hybridization techniques, and operability at low bias levels.
This all needs to be achieved while still allowing for modification of diode structure and pitch to achieve the necessary diode height to interact with the full spill of light from the Pandora grating and achieve high enough spectral resolution for accurate capture of formaldehyde transmission signals. To that end, mature 1D arrays of P-on-N photodiodes with electrical contacts on their upper surface that can fan out to wire bond pads matched to external readout electronics were selected (Figure 5(a)).
N+ and N- implant variations are being evaluated to balance depletion region depth, and therefore responsivity, with serial resistance for optimized performance. Within these constraints, diode arrays with lateral pitches of 25 µm (Figure 5(b) i) and 50 µm (Figure 5(b) ii) and heights of 200 µm to 500 µm are being fabricated to best balance the total active area ($\sim$75active area for 25 µm pitch devices due to the surface metallization cover, $\sim$ 85% active area for 50 µm pitch devices) and spatial resolution during final applications in the field.
To speed delivery, the ECI effort has relied on fully passive photodiode arrays with each diode fanning out to a wirebond output pad, which can then be connected to the wirebond input pad of a corresponding ROIC. In this case, a set of four commercial CTIA CMOS readouts with wire bondable inputs is used Figure 5(a)). ROICs will be situated on a basic PCB with output-boosting operational amplifiers to allow operation by an external commercial imaging array controller.
While this is a simple approach to readout, it is an ideal starting point for developing readouts for more complex SiC arrays, as SiC can be read with biases consistent with those already matured for silicon detectors ($\leq 2$ V) with sufficient responsivity to fill detector wells in a reasonable exposure time. This includes generalization to reading out 2D SiC arrays with similar top-side fanout structures, as will be tested in the PReSSiC effort.
@@
5 Current Testing of Photodetector Properties and Near Term Future Work
Our team has conducted initial characterization and testing of diodes during the development and fabrication process. Of note is that since much of the work is still in early stages, these photodiode characteristics are not yet optimized and in some cases bounds that do not capture the expected or even current performance of photodiodes due to evaluation techniques or hardware. This section describes some of the tests that have been conducted already and notes their limitations where applicable.
In order to achieve high-quality Formaldehyde trace signature retrieval using the Pandora platform for the ECI effort, individual photodiodes in the arrays described previously must meet minimum performance requirements. In particular, a minimum quantum efficiency (QE) of 0.14 at 300 nm is necessary to accumulate sufficient signal within a reasonable acquisition time for the selected ROIC, and dark current must be below 20 pA to achieve an appropriate signal-to-noise ratio.
Preliminary testing on isolated large photodiodes (0.505 mm2) with implant structures similar to those targeted for the ECI effort (with different profiles) meet or exceed these requirements, achieving a QE of 0.1 at 345 nm, with QE increasing as the wavelength shortens (Figure 6(a), measurements performed at 0 V bias). Performance is even more promising in terms of dark current. Extracting dark noise from a long-time measurement of photodiode current under dark conditions (Figure 6(b), data collected at 0 V bias) establishes a ceiling on dark current of approximately 6 fA, without any cooling or additional stabilization. This performance should allow for signal-to-noise ratios in excess of 1172 during measurement.
Moving forward, analysis of dark current and QE will be expanded by probing detectors coupled to the selected ROIC with integration capability, allowing for more accurate quantification of dark current. This step is essential for accurately assessing the utility of SiC photodiodes in astrophysics applications.

### v2

@@ 3 Overview of development activities
SiC instrument development by our team under the current programs have leveraged the expertise of scientists, engineers and commercial companies with fabrication, integration and implementation experience. Key to the fabrication tasks across all the project is the work that CoolCAD Electronics, LLC, has been carrying using unique expertise in the design and fabrication of devices and integrated circuits with SiC. CoolCAD has previously worked on SiC low-power CMOS, deep-UV sensors, and building blocks for deep-UV-sensitive focal plane arrays as well as high-power, high-voltage and high-temperature SiC devices. CoolCAD’s current work to produce UV sensors and readout circuitry on the same substrate leverages background work supported by NASA SBIR programs (NNX17CG32P, “Silicon Carbide 10µm-Pitch UV Imaging Array and APD with Active Pixel Readout,” Phase 1 SBIR; NNX16CG51P, “Single Chip EUV, VUV, and Deep UV Photodetector System with Integrated Amplifier,” Phase 1 SBIR; NNX17CG15C, “A Silicon Carbide Foundry for NASA’s UV and High Temperature CMOS Electronics Needs,” Phase 2/2E SBIR). CoolCAD has demonstrated two classes of 4H-SiC UV sensor devices and structures, both of which are top-lit: PN-junction and Schottky diodes, both as single diodes and in Focal Plane Arrays, which are the focus of the current projects. The spectral responsivity of a given SiC diode is, among other factors, dependent on the location and depth of the photocurrent-collecting junction electric field. Schottky diodes, with the built-in field closest to the surface, are expected to show a higher responsivity at shorter wavelengths, which has been observed in fabricated devices. In PN junction devices which are produced via implanting dopants at specific levels and concentration in a SiC lattice, CoolCAD employs implant profile engineering, to set the implant energies and doses necessary to achieve a desired profile, which has also been incorporated in current work. Finally, both past GOES-R funding and a current IRAD has supported demonstration of the radiation hardness and visible blind nature of the SiC material that is relevant to all of the projects under development.
@@
The PReSSiC project aims to deliver a lab-tested 1×128 pixel 4H-SiC detector suitable for spectrometers in miniaturized platforms that can obtain radiance observations with a higher detectivity than previous generations of NUV spectrometers from 200-340nm at a Rsimilar-$\sim$200. Key development for this project has been testing a number of design trades and optimizing design and fabrication for this goal. The project has begun with a focus on design and layout of individual sensors and small arrays. Potential single-pixel geometries and fabrication parameters are being evaluated to optimize device characteristics such as responsivity and response speed, which will be examined in the context of the science requirements. A key area of exploration is a test of implant depth, as it is valuable to create an implant profile which supports enhanced sensitivity in the spectral region of interest. There are limits on this process, such as difficulty of getting implants past a certain depth into the tough SiC crystal material. An additional consideration is that in a focal plane array, the pixels must be electrically and optically isolated by deep trenches for clear image formation. CoolCAD Electronics has previously demonstrated small-pixel-count 2-D arrays with 10 µm pitch, that had not featured deep trench isolation but is examining deep isolation trenches in the new arrays.
@@
PReSSiC is fabricating a number of different test structures and small-pixel-count arrays to identify potential fabrication difficulties and limits to our fabrication techniques. Diode pitch is a key parameter for FPA performance and is being varied between 6-40 µm to establish factors limiting diode miniaturization. In addition to pitch, PReSSiC is looking to optimize other factors influencing SiC photodiode performance including: contact resistance to minimize current losses at interfaces, minimum trench width to allow for effective fill and planarization which maximizes array density, patterning long metal lines for effective electrical interconnection lines, surface leakage reduction to improve noise performance, and identification of additional factors affecting responsivity and dark current. The team has also started examining packaging and input/output (I/O) structure design for both active and passive pixels. The team further is designing and building a set of readout electronics, particularly for planned scaling up of the pixel count for the full array. Radiometric testing is being conducted and will be conducted at numerous steps along the way using a testbed that is currently under assembly so that the arrays can eventually be tested under both ambient and vacuum conditions. Key performance testing related to responsivity, noise performance and dispersion will be compared to requirements based upon simulations.
@@
Another key ongoing development effort is the NASA Early Career Initiative-supported “Design of a Broad Utility, Visible-Blind, SiC, UV Spectrometer for Astrophysics, Planetary and Earth Science Applications, Validated Through Representative Measurement of Climate Forcing Gases” project, hereafter referred to as the Pandora+SiC effort. The Pandora+SiC effort aims to develop and integrate visible blind UV detectors into an existing, ground-based, atmospheric spectroscopy platform, Pandora, [15] which has been developed by a commercial partner, SciGlob, LLC. Pandora+SiC will improve the reliability and temporal resolution of total column formaldehyde measurements conducted by Pandora either for direct data collection or validation for data collected from satellite platforms like the ozone monitoring instrument (OMI) [32]. A simultaneous goal of this project is also to simplify instrument setup through elimination of filter components, enabled by the visible blindness of SiC, and reducing or eliminating cooling requirements, enabled by the reduced dark current of SiC. Pandora+SiC is at the point of integrating fabricated SiC detectors, $1\times 128$ and $1\times 1024$ arrays of high aspect ratio ($25~{}\mu\text{m}\times 200~{}\mu\text{m}$, $25~{}\mu\text{m}\times 500~{}\mu\text{m}$, and $50~{}\mu\text{m}\times 500~{}\mu\text{m}$), passive, P-on-N photodiodes with external capacitive transimpedance amplifier (CTIA) read-out integrated circuits (ROICs) configured to be controlled by an external imaging array controller. Following integration, detector units will be tested to confirm quantum efficiency (QE) greater than 0.1 at 345 nm and dark current $<20~{}\text{pA}$ (individual diodes probed so far show a dark current upper limit of $<6~{}\text{fA}$, discussed in section 5), and integrated into the Pandora spectrometer.
@@
The current detector and electronics configuration for both the PReSSiC detectors and the Pandora+SiC detectors has been tailored for easy integration into planetary science instruments and Pandora instrument systems, respectively. While not directly applicable to HWO, testing of pixel design, properties and layout in the PReSSiC project and maturation of the integration scheme developed through the Pandora+SiC project both provide key techology and engineering advancement for potential HWO detectors. Achieving desired noise characteristics, responsivity, and validating operation of SiC detectors using standard read out techniques offers a compelling platform for operation of denser and higher dimensionality SiC photodiode arrays of interest for use in potential Coronagraph, Spectrograph, and High-Resolution Imaging Instruments.
@@ 4 Fabrication and Integration Progress of Photodetectors
For the PReSSiC and Pandora+SiC projects, many of the initial tasks are focused on design and fabrication, and CoolCAD is working on a variety of linear photodiode structures in order to map the design space, identify design best practices, and optimize fabrication techniques.
Figure 1: Micrographs from the process of different diode regions defined for subsequent ion implantation of the tall, narrow-aspect-ratio diodes being developed for spectroscopy applications for the Pandora+SiC project.
Diode geometry is the first aspect being explored. For the particular trace gas spectroscopy application targeted by the Pandora+SiC project, the pixel pitch required for the target spectral resolution is 25 $\mu$m. The optical design of the instrument can take advantage of pixels with high aspect ratio. Therefore, the initial fabrication run of the program includes pixels with 25 $\mu$m x 500 $\mu$m and 25 $\mu$m x 200 $\mu$m drawn areas, along with test structures with a different aspect ratio. Figure 1 shows micrographs from the fabrication process of these designs, defining different diode regions. In the context of the PReSSiC program, a major program target is miniaturization to a very small pixel pitch and the devices being fabricated for this program include diodes with pitches as small as 10 $\mu$m and 6 µm. Figure 2 shows shows fabrication micrographs of some of these miniaturized structures including dopant implantation sites on a wafer which features diode arrays with 6 $\mu$m, 10 $\mu$m, 15 $\mu$m, and 25 $\mu$m pitches. Other geometric features being explored are the placement and sizes of contacts and interconnects, to evaluate certain trade-off aspects such as fill factor vs. diode series resistance.
@@
Another aspect under exploration is the utility of different junction structures. Since shorter wavelengths have larger absorption coefficients and smaller absorption depths, the different junction structures, which define the position and depth of the depletion region where photogenerated charge carriers are collected, are expected to have somewhat different spectral responsivity. To focus on broadband responsivity pushing toward 350 nm, the PReSSiC and Pandora+SiC programs are fabricating N-/P and N+/P junction diodes with different doping depths and concentrations, yielding different geometries and depths for the photocurrent collection regions. This approach shifts both the peak wavelength response as well as the bandwidth of total response, allowing tuning for specific measurements. To optimize response at shorter wavelengths for a separate NASA SBIR program, we are also fabricating Schottky diodes, which have the photocurrent collection region starting directly at the upper surface of the diode where shorter wavelengths are readily absorbed. Planned further developments include the use of different epitaxial structures, pr-dopant implantation or metallization for Schottky junction formation, to allow for more sophisticated diode fabrication as well.
Figure 3: The wafer-scale layouts for the PReSSiC project fabrication run (left) and the Pandora+SiC program fabrication run (middle, right).
Beyond individual pixel size and responsivity, in applications which use diode arrays, such as spectroscopy or imaging, sufficient pixel-to-pixel isolation is required to allow clean data collection. For the Pandora+SiC program, deep trenches between the photoactive regions of each diode are being implemented for this purpose. For the PReSSiC program, in addition to trenches, we are also implementing array variations using a specific diode geometry which was first tested for this purpose in another development program. This alternative approach allows us to increase the fill factor, especially of the smallest diodes. The development and evaluation of these isolation structures are being pursued in the SBIR program as well.
Figure 3 shows the wafer-scale designs for the initial runs in the PReSSiC project (left panel) and the Pandora+SiC program (right panel). The PReSSiC program design includes small-pixel-count arrays (15x1 and 8x1) and single test diodes. The Pandora+SiC program design includes 128x1 and 1024x1 arrays and single test diodes. Both designs are currently in fabrication and have progressed through their dopant implant and isolation trench formation steps. We are currently optimizing back-end processing steps, which comprises the metallization steps for contact metal and interconnect formation, preceding implementation of these steps on the PReSSiC and Pandora+SiC wafers to complete their fabrication.
@@
For both projects, the next stage of development will include active diode arrays, for which the first stage of a readout circuit converting the photocurrent to a voltage level over an integration period is fabricated on the same die with the photodiode arrays. CoolCAD Electronics has active development programs on SiC MOSFET structures and circuits and has previously demonstrated single diodes integrated with MOSFET baser read out circuits. Figure 4 shows previously-fabricated examples of this structure. These development programs all start from the level of physics-based semiconductor device modeling and simulations, and include process simulation and development efforts as well as device and circuit design, layout, and fabrication. For the PReSSiC project, in addition to one-dimensional arrays for spectroscopy, two-dimensional arrays with expanded pixel counts will also be implemented for push-broom spectroscopy and imaging applications. The pixel count will be expanded as well. In the SBIR program, CoolCAD is further developing larger scale two-dimensional arrays, both of passive and active pixels. For the two-dimensional active array development, this effort includes work to miniaturize the transistors so they can be implemented within a given pixel, as opposed to be connected externally with metal fan out, permitting effective scaling in 2 dimensions.
@@
Figure 5: The structure of SiC photodiode-based focal plane arrays for the Pandora+SiC effort consists of: (a) A core 1×1024 array of SiC photodiodes, which fans out to wire bond pads matched to the inputs of wire bondable ROICs. In this case, the array is interfaced with 4 CTIA ROICs, each with 256 inputs. (b) The Pandora+SiC effort is currently in the process of characterizing multiple diode variants, including: (i) arrays with a 25 µm pitch, and (ii) arrays with a 50 µm pitch.
In addition to advancing fundamental understanding of SiC photodiode fabrication and performance, the Pandora+SiC effort is focused on a short-term field demonstration of SiC focal plane arrays in an existing atmospheric spectroscopy system, Pandora, for formaldehyde detection. As a result, the selection of diode array formats was informed by the need for a mature and highly reliable diode design, the ability to interface with external readout circuitry without complex hybridization techniques, and operability at low bias levels.
This all needs to be achieved while still allowing for modification of diode structure and pitch to achieve the necessary diode height to interact with the full spill of light from the Pandora grating, in order to afford a high enough spectral resolution for accurate capture of formaldehyde transmission signals. To that end, mature 1D arrays of PN photodiodes with electrical contacts on their upper surface that can fan out to wire bond pads matched to external readout electronics were selected (Figure 5(a)).
P/N+ and P/N- implant variations are being evaluated to balance depletion region depth, and therefore responsivity, with serial resistance for optimized performance. Within these constraints, diode arrays with lateral pitches of 25 µm (Figure 5(b) i) and 50 µm (Figure 5(b) ii) and heights of 200 µm to 500 µm are being fabricated to best balance the total active area ($\sim$75active area for 25 µm pitch devices due to the surface metallization cover, $\sim$ 85% active area for 50 µm pitch devices) and spatial resolution during final applications in the field.
To speed delivery, the Pandora+SiC effort has relied on fully passive photodiode arrays with each diode fanning out to a wirebond output pad, which can then be connected to the wirebond input pad of a corresponding ROIC. In this case, a set of four commercial CTIA CMOS readouts with wire bondable inputs is used Figure 5(a)). ROICs will be situated on a basic PCB with output-boosting operational amplifiers to allow operation by an external commercial imaging array controller.
While this is a simple approach to readout, it is an ideal starting point for developing readouts for more complex SiC arrays, as SiC can be read with biases consistent with those already matured for Si detectors ($\leq 2$ V) with sufficient responsivity to fill detector wells in a reasonable exposure time. This includes generalization to reading out 2D SiC arrays with similar top-side fanout structures, as will be tested in the PReSSiC effort.
@@
5 Current Testing of Photodetector Properties and Near-Term Future Work
Our team has conducted initial characterization and testing of diodes during the development and fabrication process. Of note is that since much of the work is still in early stages, these photodiode characteristics are not yet optimized and in some cases measurements are limited by evaluation techniques or hardware. Where this is the case, we provide a minimum possible performance parameter but expect SiC photodiodes to significantly exceed values reported here. This section describes some of the tests that have been conducted already and notes their limitations where applicable.
In order to achieve high-quality formaldehyde trace signature retrieval using the Pandora platform for the Pandora+SiC effort, individual photodiodes in the arrays described previously must meet minimum performance requirements. In particular, a minimum quantum efficiency (QE) of 0.14 at 300 nm is necessary to accumulate sufficient signal within a reasonable acquisition time for the selected ROIC, and dark current must be below 20 pA to achieve an appropriate signal-to-noise ratio.
Preliminary testing on isolated large photodiodes ($\approx$ 0.5 mm2) with implant structures similar to those targeted for the Pandora+SiC effort (with different profiles) meet or exceed these requirements, achieving a QE of 0.1 at 345 nm, with QE peaking at $\approx$ 0.4-0.5 at 280 nm (Figure 6(a), measurements performed at 0 V bias). Performance is even more promising in terms of dark current. Extracting dark noise from a long-duration measurement of photodiode current under dark conditions and in the absence of any cooling or additional stabilization (Figure 6(b), data collected at 0 V bias) establishes a ceiling on dark current of approximately 6 fA. We claim this as a ceiling as this value is likely a result of shot noise arising from the transresistance amplifier used for the measurement.. This performance should conservatively allow for signal-to-noise ratios in excess of 1172 during measurement.
Moving forward, analysis of dark current and QE will be expanded by probing detectors coupled to the selected CTIA ROIC with integration capability and correlated double sampling, allowing for more accurate quantification of dark current removing the effects of shot noise. This step is essential for accurately assessing the utility of SiC photodiodes in astrophysics applications.